OpenStudy (anonymous):

what is the equation of the chord with the given mid pt (h,k) in the rectangular hyperbola of form x y=c^2

OpenStudy (swag):

.

OpenStudy (shubhamsrg):

well i dont know much about hyperbolas but still maybe i can manage to attempt this question : let one point of the chord on the hyperbola be (t, c^2 /t ) so other point = (2h -t , 2k - c^2 /t) => ( 2h-t) ( 2k - c^2 /t) = c^2 on simplification, it'll give a quadratic eqn in t .. so we can solve for t easily and thus for the eqn of the line.. hope i could help..